OpenStudy (anonymous):

for questions 9-10 what are the solutions of the equation A. c^2-4c=0 B. z^2-4c-27=0

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

\[\huge c^2- 4c=0 \rightarrow c(c-4)= 0\] \[\huge \rightarrow c =0 or (c-4)= 0\] \[\huge \rightarrow c =0 or c=4\] c= 0, 4

OpenStudy (anonymous):

ok thanks what about B

OpenStudy (anonymous):

\[\huge z^2-4c-27 =0\] Let us put c=0 , we find; \[\huge z^2-4\times 0-27 =0 \rightarrow z^2- 0-27 =0\] \[\huge \rightarrow z^2- 27 =0 \rightarrow z^2 = 27 \rightarrow z= \pm 3\]

OpenStudy (anonymous):

i goofed up its z^2-4z-27=0 does that change any thing

OpenStudy (anonymous):

At c = 4, we find \[\huge \rightarrow z^2- 4 \times 4-27 =0\] \[\huge \rightarrow z^2- 16-27 =0\] \[\huge \rightarrow z^2- 16-27 =0 \rightarrow z^2- 43=0\] \[\huge \rightarrow z^2= 43 \rightarrow z= \pm \sqrt {43}\]

OpenStudy (anonymous):

@handyandy3

OpenStudy (anonymous):

my answer choices are 3,9 3,-9 -9,3 -3,-9

OpenStudy (anonymous):

@handyandy3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so are you able to figure it out?

OpenStudy (anonymous):

@dpasingh